OpenStudy (anonymous):

Can someone help me solve this: -2x - 15 = 6x + 9

OpenStudy (dragons1415):

do you need all the steps?

OpenStudy (dragons1415):

bcuz the answer turns out to be x=-3

OpenStudy (anonymous):

These are my options: 3 -3 -1 6 And yes, I need to know how to do it

OpenStudy (ajprincess):

Bring the terms with x to one side and number to the other side. -2x-15=6x+9 Add 2x both sides -2x-15+2x=6x+9+2x -15=6x+9+2x Subtract 9 both sides -15-9=6x+9+2x-9 -15-9=6x+2x Nw solve for x. Does that help?

OpenStudy (anonymous):

YES! Thank you both so much! ^-^

OpenStudy (ajprincess):

Welcome:)